OpenStudy (anonymous):

Given that Sin 0=3/5, find the values of the other five triganomic functions

OpenStudy (jmark):

Given Sin θ=3/5 =>we have sin^2θ+cos^2θ=1 =>(3/5)^2+cos^2θ=1 =>cos^2θ=1-(9/25) =>cos^2θ=16/25 =>cos θ=4/5 we have tan θ=sin θ/cos θ =>tan θ=(3/5)/(4/5) =>tan θ=3/4 =>cot θ=1/tan θ =>cot θ=1/(3/4) =>cot θ=4/3 =>cosec θ= 1/sin θ =>cosec θ=1/(3/5) =>cosec θ=5/3 =>sec θ=1/cos θ =>sec θ=1/(4/5) =>sec θ=5/4 sample paper for class 12 CBSE 2013 http://goo.gl/RYLqP6